# A Spatial-statistical model to analyse historical rutting data

### 08 Jan 2024 v1

@@ 4.2.1 Investigating areas with high rutting rates
Figure 6: Graph (a): Expected rutting inferred from the spatial rutting component $\hat{\omega}(s)$ plus the ‘best’ (or ‘worse’) explanatory values at road segments $2100-2500$ towards Sweden. The estimate $\hat{\eta}_{\mathrm{Agc}}$ is for pavement type Agc plus the random yearly effect and the spatial rutting component. Graph (b): expected rutting inferred from $\hat{\omega}(s)$ plus the random yearly effect $\hat{\gamma}$ and each explanatory variable added sequentially in the model. Graph (c): bearing capacity for year 2021 (tonn), with 16 tonn or greater indicating a good parameter. Graphs (d) and (e) give the ratios of surface curvature index (SCI) to base curvature index (BCI) for road segments 2100-2500 (metres) or the whole E14, respectively, with values below $12$ indicating a good curvature index.
@@
The above results suggest pavement type as the driving explanatory variable for expected rutting (Figure 6), so we determine its importance on rutting by changing the asphalt type from the true one (i.e., asphalt gravel concrete, Agc) to the stronger asphalt type Sma (stone mastic asphalt) and keeping all other variables (e.g., lane width) as given in the data set. The remaining explanatory variables are then added sequentially in the model with pavement type, i.e., stacking. Pavement type (Agc or Sma) is first added to the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$. We denote this by: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$, respectively. Next, rut depth from the previous year is added, and it is represented as $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Sma(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$. Similarly, the model with lane width $\{\hat{\beta}_{w}z_{\mathrm{w(s,t)}}\}$ added is represented by $\hat{\tilde{\eta}}_{\mathrm{w}}$. The results are shown in Figure 6, graph (b).
Results from using stone mastic asphalt (Sma) are represented by the solid lines, and from using asphalt gravel concrete (Agc), the dotted lines. The results confirm that pavement type is the driving factor of rutting. If the stronger binder type Sma were laid at this stretch instead of the true pavement type Agc, expected rutting would decrease, reaching at most $4.35$ mm/year (compared with 4.68 mm/year). This is approximately 0.33 mm/year less rutting annually, with an expected lifetime of 6 years, i.e., 1 extra year.
@@
We also look at the available data for bearing capacity (Figure 6, graph (c)), surface curvature index (SCI) and base curvature index (BCI) for road segments 2100-2500. The data is available for year 2021, but it is a good indicator of the construction quality and the soil to sustain pressure over the past decade; Figure 6, graph (c). For bearing capacity, values greater than $16$ tonns (       ) suggest a good foundation design. The ratio of SCI to BCI ((Figure 6, graphs (d) and (e)) gives an indication of the location of the deflection (deformation) in the pavement layer (surface or base). The data suggests good bearing capacity, and that high rutting may be a result of other influential factors, e.g., side slope height (data not available), lack of side ditches to effect water drainage or deformation in the sub-base layer developed in thawing periods. We note that the bearing capacity data was sampled during the summer and corrected for temperature. During the summer, the measurements can be higher due to dry conditions, compared to those in thawing periods with high moisture content. Permanent deformation could have occurred during thawing periods, resulting in high rutting. There is also the possibility of varying construction material quality at various parts of the road, even though the pavement type (asphalt gravel concrete) is the same for the entire stretch (road segments $1500-3355$). Therefore, we investigate the consistency in the pavement construction by looking at deflections in the surface layer and the subsoil or reinforcement layer. The ratio of SCI to BCI is a good indication of the strengths or weaknesses in the construction, with $\mathrm{SCI/BCI}<12$ (       ) indicating a good parameter value. If this ratio is constant, then the construction is considered even, otherwise, there are variations indicating deformation in the various layers. The data, suggests an uneven construction of E14, with deeper deflections in the base layer.
We conclude that pavement type is the most important contributing factor to rutting. So, stronger binder types are recommended to ensure longer expected lifetimes.
4.3 Estimated model parameters
The estimated parameters and 95% credible intervals (CI) for the spatial Model 1 and non-spatial Model 4 are reported in Table 3. Comparing the coefficients for the explanatory variables between spatial models and their non-spatial version, we observe that the spatial models have considerably larger estimates for the regression coefficients of pavement type (see also Figure A.2 in Appendix A). Further, we see that the variance of the ‘noise’ term ($\hat{\sigma}_{\varepsilon}$) is larger for the models without the spatial rutting component. One explanation for this result is that the spatial rutting component account for unobservable factors, e.g., weather that vary in space, the effect of studded tires and other influencing factors such as traffic type that are not included in the model. Varying weather patterns across Norway and towards the Swedish border will affect the function of surface layers of pavement construction. In the winter time, increased temperatures result in a shorter time of frozen ground and snow on the surface and changes in the number of freeze-thaw cycles, thus, negatively affecting lower-grade binder types such as asphalt gravel concrete (Agc) and leading to increasing rutting. Asphalt gravel concrete is laid from () $(\sim)\ 30$ kilometres towards the Swedish border (Figure 1 (a)); an area that lie in the sub-artic climate zone with long severe winters and short warm to cool summers Peel et al. (2007), Ketzler et al. (2021), and having poorly developed networks (rural E14 highway west of Östersund) Jenelius (2010). This area is also more likely to be traversed by heavy-traffic (exceeding $3500$ kilograms or $3.5$ tonnes) Hovi et al. (2018), Pinchasik et al. (2020), negatively affecting pavement performance. More of the expected rutting is also explained by rut depth and lane width from the spatial model compared with the non-spatial model. This can be seen by the smaller values of the ‘noise‘ terms ($\hat{\sigma}_{\varepsilon}$) and uncertainty in the random year effect ($\hat{\sigma}_{\gamma}$) from the spatial model (see Table 3 and Table 4 in Appendix A for the simpler models).
Table 3: Estimated model parameters with $95\%$ credible intervals
Spatial Model Non-spatial Model
Model 1 Model 4
Mean/Median 95% CI Mean/Median 95% CI
$\hat{\beta}_{\mathrm{Ac}}$ 7.13 [6.62, 7.65] 3.25 [2.81, 3.68]
$\hat{\beta}_{\mathrm{Agc}}$ 7.34 [6.92, 7.76] 3.21 [2.75, 3.67]
$\hat{\beta}_{\mathrm{Sma}}$ 3.63 [3.37, 3.90] 2.00 [1.81, 2.19]
$\hat{\beta}_{\mathrm{d-1}}$ 0.74 [0.66, 0.82] 1.08 [0.98, 1.18]
$\hat{\beta}_{\mathrm{w}}$0.10 [-0.14, -0.07] 0.05 [-0.02, 0.11]
$\hat{\sigma}_{\omega(s)}$ 0.47 [0.44, 0.51] - -
$\hat{\sigma}_{\gamma}$ 0.72 [0.62, 0.88] 1.38 [0.96, 2.16]
$\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.54 [1.53, 1.56]
4.3.1 Effects of the explanatory variables
Interacting with the AADT, which is spatially varying, the effect of asphalt concrete ($\hat{\beta}_{\mathrm{Ac}}$) and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) on rutting may be explained by: heavy-traffic ($>3.5$ tonnes), whose proportion changes along the road and poor network design as these are constant in time. Also, various parts of the pavement have several layers due to non-removal during maintenance, leading to variability in pavement performance. The uncertainty from asphalt concrete and asphalt gravel concrete ($\hat{\beta}_{\mathrm{Agc}}$) are also larger (wider 95% credible interval) compared with stone mastic asphalt ($\hat{\beta}_{\mathrm{Sma}}$), but accounting for spatial dependencies allows more of the variance to be explained. Rut depth from the previous year ($\hat{\beta}_{\mathrm{d-1}}$) becomes more important when the spatial variability in pavement rutting is not considered (Model 4, Table 3). Further, instead of explaining the variability in expected rutting by pavement type, the non-spatial models explain this variability between years (Table 3 and Figure 3 (b)). The estimated effect of lane width ($\hat{\beta}_{\mathrm{w}}$) is small and negative. This suggests that narrower lane widths would give higher expected rutting.
Based on these results, we conclude that changes in the coefficients are large when spatial dependencies of factors that are not modelled because of non-observable data (e.g., weather) are considered. And, the uncertainty explained is greater (smaller $\hat{\sigma}_{\varepsilon}$) for the spatial model.
@@ 5 How road authorities can make use of this model
The spatial-statistical model serves as a potent tool for road authorities, offering a data-driven approach to enhance maintenance strategies, identify critical areas for intervention and optimize road infrastructure durability. This approach facilitates a more ultimate diagnosis of rutting, empowering authorities to make informed decisions for sustainable road infrastructure.
Road authorities can use the spatial-statistical model for historical rutting data in several ways to improve maintenance strategies and prioritize interventions. These include; i) the identification of ‘hot spots’ — the spatial model helps pinpoint hot spots where rutting is accelerating, indicating areas in urgent need of maintenance. By analyzing coefficients of factors known to influence rutting (e.g., asphalt type), random effects, and spatial components, authorities can prioritize resources for targeted interventions in these critical locations; ii) comprehensive rutting diagnostic and understanding — the model provides a more comprehensive diagnostic and understanding of rutting. Explanatory variables and random yearly effects alone may not capture the full variability, but the inclusion of spatial components enhances the model’s explanatory power. This comprehensive understanding allows road authorities to make informed decisions about maintenance priorities; iii) extended expected lifetime — considering both spatial and random yearly effects extends the expected lifetime estimation of road sections. This information is valuable for long-term planning, enabling authorities to proactively address maintenance needs and allocate resources efficiently; iv) pavement type considerations — the model highlights the significant impact of pavement type on rutting rates. Road authorities can use this information to make informed decisions about pavement materials during construction or rehabilitation, optimizing durability and extending the overall lifetime of the road infrastructure. The recommendation of stronger asphalt binders, such as stone mastic asphalt (Sma), to enhance pavement durability, provides practical guidance for selecting materials that contribute to longer-lasting road surfaces; v) investigation of contributing factors — in-depth investigations into high rutting areas reveal potential contributing factors beyond pavement type, such as drainage issues, heavy traffic, and road design flaws. Road authorities can address these factors to prevent or mitigate rutting, ensuring a more holistic approach to road maintenance; vi) spatial model advantages — comparisons with non-spatial models highlight the advantages of including spatial components. The larger estimates for pavement type coefficients in spatial models emphasize the importance of spatial factors in explaining rutting variations. This information reinforces the importance of using spatial models for accurate and reliable inference and predictions; and finally vii) data-driven decision-making — the estimated model parameters and credible intervals provide a robust foundation for data-driven decision-making. Road authorities can rely on these parameters to make informed choices regarding maintenance priorities, resource allocation, and infrastructure improvements.
@@ 6 Conclusion
This research provides valuable insights into the factors influencing pavement rutting and the measurement of rut depth. The study confirms that the selection of asphalt type significantly affects rutting performance, with rut-resistant mixtures proving to be more effective in mitigating rutting. Lane width also plays a role in pavement rutting. Narrower roads tend to experience more concentrated traffic loads, which can lead to increased rutting. Wider roads, on the other hand, distribute the traffic load over a larger area, reducing the likelihood of rut formation. The data analysis indicated that narrower road stretches experienced higher rut depths. This outcome can be attributed to increased traffic volumes and higher wheel loads on roads, leading to more significant rutting. The findings, therefore, stress the importance of considering lane width when designing and maintaining pavements to minimize rutting distress. The recommendation is to incorporate a life-cycle cost-benefit perspective in the road design and planning process to minimize the potential for rutting issues over time. Further research is encouraged to explore additional factors influencing rutting; some of which include bearing capacity (Hopkins et al. 1998) as inadequate bearing strength of soil subgrades will not withstand construction traffic loading; ditch depth information that allows outflow of water from pavement structure; and lane width. This would likely increase the credibility of research results and the applicability of models to other road networks. In addition, it would be worthwhile to evaluate the long-term performance of rut-resistant asphalt mixtures in various traffic and environmental conditions.
@@ Acknowledgement
This work was supported by the Norwegian University of Science and Technology in collaboration with Statens Vegvesen Project [grant number 9056608 – SMARTere Vedlikehold (2019-2023)].
@@ Appendix A Inference from model parameters
SPATIAL MODELS Model 1 Model 2 Model 3 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 7.13 [6.62, 7.65] 7.19 [6.66, 7.72] 6.98 [6.44, 7.53] $\hat{\beta}_{\mathrm{Agc}}$ 7.34 [6.92, 7.76] 7.38 [6.95, 7.81] 7.15 [6.72, 7.59] $\hat{\beta}_{\mathrm{Sma}}$ 3.63 [3.37, 3.90] 3.57 [3.30 3.85] 3.36 [3.08, 3.64] $\hat{\beta}_{\mathrm{d-1}}$ 0.74 [0.66, 0.82] 0.77 [0.69, 0.85] - - $\hat{\beta}_{\mathrm{w}}$0.10 [-0.14, -0.07] - - - $\hat{\sigma}_{\omega(s)}$ 0.47 [0.44, 0.51] 0.48 [0.45, 0.53] 0.52 [0.48, 0.57] $\hat{\sigma}_{\gamma}$ 0.72 [0.62, 0.88] 0.81 [0.64, 1.09] 0.72 [0.58, 0.94] $\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.38 [1.37, 1.39] 1.39 [1.38, 1.40] NON-SPATIAL MODELS Model 4 Model 5 Model 6 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 3.25 [2.81, 3.68] 3.25 [2.81, 3.69] 2.72 [2.28, 3.16] $\hat{\beta}_{\mathrm{Agc}}$ 3.21 [2.75, 3.67] 3.20 [2.74, 3.67] 2.71 [2.24, 3.18] $\hat{\beta}_{\mathrm{Sma}}$ 2.00 [1.81, 2.19] 2.02 [1.83, 2.21] 1.55 [1.36, 1.73] $\hat{\beta}_{\mathrm{d-1}}$ 1.08 [0.98, 1.18] 1.08 [0.98, 1.18] - - $\hat{\beta}_{\mathrm{w}}$ 0.05 [-0.02, 0.11] - - - - $\hat{\sigma}_{\gamma}$ 1.38 [0.96, 2.16] 1.45 [0.98, 2.21] 1.43 [0.99, 2.25] $\hat{\sigma}_{\varepsilon}$ 1.54 [1.53, 1.56] 1.54 [1.53, 1.56] 1.56 [1.54, 1.57]

### 14 Feb 2024 v2

@@ 4.4.1 Exploring the spatial model with various values of the explanatory variables
Figure 6: Graph (a) displays expected rutting using spatial rutting component $\hat{\omega}(s)$ for road segments $2100-2700$, along with the ’best’ or ’worst’ explanatory values. $\hat{\eta}_{\mathrm{Agc}}$ represents pavement type Agc, random yearly effect $\hat{\gamma}$, and $\hat{\omega}(s)$. Graph (b) illustrates expected rutting inferred from $\hat{\omega}(s)$, $\hat{\gamma}$, and sequentially added explanatory variables. Graph (c) shows 2021 bearing capacity in tons; values $\geq$ 16 indicate good parameters. Graphs (d) and (e) depict SCI to BCI ratios for a short stretch or the entire E14; values $<$ 12 indicate good curvature index.
@@
We examine the impact of pavement type on rutting (Figure 6, graph (b)). Changing from asphalt type Agc (asphalt gravel concrete) to the stronger Sma (stone mastic asphalt) while keeping all other variables constant, we sequentially add additional explanatory variables to the model, starting with pavement type. Initially, pavement type (Agc or Sma) is included with the random yearly effect $\hat{\gamma}$ and the spatial rutting component $\hat{\omega}(s)$ as follows: $\hat{\tilde{\eta}}_{\mathrm{Agc}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{Sma}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Agc(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$. Subsequently, the previous year’s rut depth is incorporated, represented as $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Agc}}z_{\mathrm{Agc(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$ or $\hat{\tilde{\eta}}_{\mathrm{d-1}}=\hat{\beta}_{\mathrm{Sma}}z_{\mathrm{Sma(s,t)}}+\hat{\beta}_{\mathrm{d-1}}z_{\mathrm{d-1(s,t)}}+\hat{\gamma}+\hat{\omega}(s)$. Similarly, the model with added lane width ${\hat{\beta}_{w}z_{\mathrm{w(s,t)}}}$ is denoted as $\hat{\tilde{\eta}}_{\mathrm{w}}$. Results from using stone mastic asphalt (Sma) are depicted by solid lines, and asphalt gravel concrete (Agc) results are shown with dotted lines.
The findings affirm that pavement type significantly influences rutting. Implementing the stronger binder type Sma instead of the actual pavement type Agc would decrease expected rutting, reaching a maximum of $4.35$ mm/year (compared to 4.68 mm/year), approximately reducing rutting by 0.33 mm/year annually. Over a projected lifetime of 6 years, this equates to an additional year of serviceability.
4.4.2 Evaluating pavement subgrade support on expected rutting
@@
We examine bearing capacity, surface curvature index (SCI), and base curvature index (BCI) for road segments 2100-2700 using 2021 data, indicative of construction quality and soil pressure sustainability over the past decade (Figure 6, graph (c)). Bearing capacity values exceeding 16 tons (       ) suggest a robust foundation design. The SCI to BCI ratio (Figure 6, graphs (d) and (e)) reveals deflection location within the pavement layers. High rutting may result from factors like side slope height (data unavailable), inadequate drainage, or sub-base layer deformation during thawing periods. Bearing capacity data, sampled and temperature-corrected during summer, may inflate measurements compared to thawing periods. Varying construction material quality along the road stretch (segments 1500-3355) warrants investigation into surface and subsoil layer deflections. An SCI/BCI ratio below 12 (       ) indicates robust construction. Inconsistent ratios suggest layer deformation, with data indicating uneven construction on E14, featuring deeper base layer deflections.
In conclusion, pavement type significantly influences rutting, recommending stronger binder types for extended lifespan.
@@ 5 How road authorities can make use of this model
The spatial-statistical model serves as a potent tool for road authorities, offering a data-driven approach to enhance maintenance strategies, identify critical areas for intervention, and optimize road infrastructure durability. This approach facilitates a more precise diagnosis of rutting, enabling authorities to make informed decisions for sustainable road infrastructure.
Our analyses reveal that rutting exceeding 2 mm/year is expected at several locations along E14. Given the expected lifetime of the pavement and recommended treatment levels, this level of rutting is concerning. Maintenance carried out at rut depths of 20-25 mm, depending on traffic, and pavement designs expected to last between 20-30 years indicate that some sections of the road may only survive for at most 12.5 years (10 years in high-traffic areas). It’s imperative to consider spatial rutting components; models fitted with variables known to influence rutting and random yearly effects alone underestimate rutting variability, thus inflating pavement lifetime estimates. Including the spatial rutting component reduces estimated pavement lifetimes to at most 6 years, compared to 5 years without it. This information is crucial for road authorities’ long-term planning and resource allocation.
We’ve demonstrated the spatial-statistical model’s ability to identify areas with unexpectedly high rutting, pinpointing hot spots where rutting is accelerating and indicating urgent maintenance needs. By analyzing factors influencing rutting, random effects, and spatial components, authorities can prioritize resources for targeted interventions. Additionally, the model provides a comprehensive understanding of rutting, enhancing authorities’ ability to make informed decisions about maintenance priorities and pavement materials during construction or rehabilitation.
The model can investigate contributing factors to rutting beyond pavement type, such as drainage issues, heavy traffic, and road design flaws. Comparisons with non-spatial models emphasize the importance of spatial factors in explaining rutting variations, reinforcing the need for spatial models for accurate inference and predictions. The estimated model parameters and credible intervals provide a robust foundation for data-driven decision-making, enabling road authorities to make informed choices regarding maintenance priorities and infrastructure improvements.
@@ 6 Conclusion
This research offers valuable insights into the rutting phenomenon by elucidating both well-established factors and those that remain unexplained. It shows the importance of incorporating spatial information to accurately estimate rutting variability, showing that observations in close proximity exhibit higher similarity than those at a distance. To achieve a comprehensive understanding, it is imperative to consider both observable elements like traffic and temperature, as well as unobservable factors like weather, and integrate spatial rutting components.
Moreover, the study confirms the significant impact of asphalt type selection on rutting performance, with rut-resistant mixtures proving more effective in mitigating rutting. Lane width also influences pavement rutting, with narrower roads experiencing more concentrated traffic loads and, consequently, increased rutting. Conversely, wider roads distribute traffic loads over a larger area, reducing rut formation. The analysis reveals that wider road sections tend to experience higher rut depths compared to narrower road stretches. This result can be attributed to the increased traffic volumes and higher wheel loads present on wider roads, which contribute to more pronounced rutting. To address this issue and minimize rutting problems over time, it is advisable to integrate a life-cycle cost-benefit perspective into road design and planning processes.
Further research is encouraged to explore additional factors influencing rutting, such as bearing capacity, ditch depth for water outflow, and lane width. This would enhance the credibility of research results and the applicability of models to other road networks. Additionally, evaluating the long-term performance of rut-resistant asphalt mixtures under various traffic and environmental conditions would be worthwhile.
@@ Acknowledgement
This work was supported by the Norwegian University of Science and Technology in collaboration with Statens Vegvesen Project [grant number 9056608 – Smarter Maintenance (2019-2023)].
@@ Appendix A Inference from model parameters
SPATIAL MODELS Model 1 Model 2 Model 3 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 7.09 [6.56, 7.61] 7.08 [6.55, 7.61] 6.89 [6.34, 7.44] $\hat{\beta}_{\mathrm{Agc}}$ 7.31 [6.88, 7.73] 7.29 [6.86, 7.71] 7.07 [6.64, 7.50] $\hat{\beta}_{\mathrm{Sma}}$ 3.53 [3.26, 3.80] 3.55 [3.28 3.82] 3.31 [3.03, 3.59] $\hat{\beta}_{\mathrm{d-1}}$ 0.76 [0.68, 0.84] 0.76 [0.68, 0.84] - - $\hat{\beta}_{\mathrm{w}}$ 0.07 [0.01, 0.13] - - - $\hat{\sigma}_{\omega(s)}$ 0.48 [0.45, 0.52] 0.48 [0.45, 0.52] 0.52 [0.48, 0.57] $\hat{\sigma}_{\gamma}$ 0.92 [0.63, 1.42] 0.93 [0.63, 1.40] 0.93 [0.63, 1.42] $\hat{\sigma}_{\varepsilon}$ 1.38 [1.38, 1.39] 1.38 [1.38, 1.39] 1.39 [1.38, 1.40] NON-SPATIAL MODELS Model 4 Model 5 Model 6 Mean/Median 95% CI Mean/Median 95% CI Mean/Median 95% CI $\hat{\beta}_{\mathrm{Ac}}$ 3.27 [2.84, 3.71] 3.28 [2.85, 3.72] 2.72 [2.29, 3.16] $\hat{\beta}_{\mathrm{Agc}}$ 3.24 [2.78, 3.70] 3.23 [2.77, 3.70] 2.71 [2.25, 3.17] $\hat{\beta}_{\mathrm{Sma}}$ 2.03 [1.84, 2.22] 2.05 [1.86, 2.24] 1.55 [1.36, 1.73] $\hat{\beta}_{\mathrm{d-1}}$ 1.07 [0.98, 1.17] 1.08 [0.98, 1.18] - - $\hat{\beta}_{\mathrm{w}}$ 0.05 [ -0.02, 0.11] - - - - $\hat{\sigma}_{\gamma}$ 1.37 [0.94, 2.10] 1.37 [0.94, 2.10] 1.41 [0.96, 2.17] $\hat{\sigma}_{\varepsilon}$ 1.54 [1.53, 1.56] 1.56 [1.54, 1.57] 1.56 [1.54, 1.57]